OpenStudy (anonymous):

Which shows the expression below simplified? 0.17 + (3.5 × 10-2)

OpenStudy (anonymous):

3.67 × 10-2 2.05 × 10-2 3.67 × 10-3 2.05 × 10-1

OpenStudy (anonymous):

is that plus or times?

OpenStudy (anonymous):

Its plus then times.

OpenStudy (anonymous):

I got C?

OpenStudy (anonymous):

or if you prefer, \(2.05\times 10^{-1}\)

OpenStudy (anonymous):

no not C, try D

OpenStudy (anonymous):

Lets see

OpenStudy (anonymous):

19.5?

OpenStudy (anonymous):

If your helping someone else you can help them and take your time and come back whenever. (:

OpenStudy (anonymous):

\[3.5\times 10^{-2}=0.035\] and \[.035+.17=.205\]

OpenStudy (anonymous):

in scientific notation \[.205=2.05\times 10^{-1}\]

OpenStudy (anonymous):

i think that is choice D

OpenStudy (anonymous):

Thank you let me check!

OpenStudy (anonymous):

ok you check

OpenStudy (anonymous):

You were right! Thank you so much, someday we should talk to each other

OpenStudy (anonymous):

sure. y ou got other questions or is that it?

OpenStudy (anonymous):

I got others

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

k post!